# CAKE: Cascading and Adaptive KV Cache Eviction with Layer Preferences

### v1

@@ Abstract
Large language models (LLMs) excel at processing long sequences, boosting demand for key-value (KV) caching. While recent efforts to evict KV cache have alleviated the inference burden, they often fail to allocate resources rationally across layers with different attention patterns. In this paper, we introduce Cascading and Adaptive KV cache Eviction (CAKE), a novel approach that frames KV cache eviction as a “cake-slicing problem.” CAKE assesses layer-specific preferences by considering attention dynamics in both spatial and temporal dimensions, allocates rational cache size for layers accordingly, and manages memory constraints in a cascading manner. This approach enables a global view of cache allocation, adaptively distributing resources across diverse attention mechanisms while maintaining memory budgets. CAKE also employs a new eviction indicator that considers the shifting importance of tokens over time, addressing limitations in existing methods that overlook temporal dynamics. Comprehensive experiments on LongBench and NeedleBench show that CAKE maintains model performance with only 3.2% of the KV cache and consistently outperforms current baselines across various models and memory constraints, particularly in low-memory settings. Additionally, CAKE achieves over 10$\times$ speedup in decoding latency compared to full cache when processing contexts of 128K tokens with FlashAttention-2. Our code is available at https://github.com/antgroup/cakekv.
@@ 1 Introduction
Large language models (LLMs) (Zhao et al., 2023; Achiam et al., 2023; Dubey et al., 2024; Anthropic, 2024; AI, 2024) have enhanced their long-text processing capabilities, improving performance in multi-turn dialogues (Chiang et al., 2023), document summarization (Zhang et al., 2024a), question answering (Kamalloo et al., 2023), and information retrieval (Liu et al., 2024c). New models such as GPT-4 (Achiam et al., 2023), Claude 3.5 (Anthropic, 2024), LLaMA 3.1 (Dubey et al., 2024) and Mistral Large 2 (AI, 2024) have extended token processing capacities beyond 128K. Expanded contexts necessitate a linear increase in key-value (KV) cache size, resulting in heavier inference-time memory burdens. Shazeer (2019); Ainslie et al. (2023) partially address this issue by merging key-value heads during the training phase. However, optimizing key-value cache without additional training is crucial for efficient inference of long contexts under memory constraints, particularly in typical deployment scenarios where the model structure is fixed.
One way to maintain a manageable KV cache size on the fly is to remove some KV pairs (Xiao et al., 2023; Zhang et al., 2024b; Li et al., 2024b). The idea is to eliminate less important KV pairs based on certain rules. Although recent methods have enhanced pair selection for removal, they typically assign uniform cache sizes across layers, disregarding layer-specific requirements. This approach can impair performance under memory constraints. Recent research is trying to improve this by looking at how attention works in different layers (Yang et al., 2024a; Cai et al., 2024). These methods, which depend on prior observations, might not work well for all input types and models. Another idea is to adjust the cache size based on the current attention pattern (Wan et al., 2024). This can help, but it may not always optimize overall performance.
@@
Allocating optimal cache sizes for different layers with a fixed memory budget is akin to “cake-slicing.” The crux of this challenge is assessing each layer’s affinity for KV cache, which is mirrored in its attention mechanisms. Our research delves into these mechanisms, examining spatial and temporal aspects across layers. Spatially, attention distribution across tokens can differ markedly between layers. Some layers may spread attention broadly, while others concentrate on specific tokens, as depicted in Figure 1(a) and Figure 1(b). Temporally, attention hotspots in certain layers shift over time in Figure 1(c), whereas in others, they remain constant in Figure 1(d). The variability of attention pattern necessitates a dynamic approach to memory allocation that goes beyond uniform (Zhang et al., 2024b; Liu et al., 2024d; Ren & Zhu, 2024; Li et al., 2024b), depicted in Figure 2(a), or fixed-pattern methods (Cai et al., 2024), drawn in Figure 2(b). They lack the flexibility to adapt to the nuanced attention dynamics. To enhance cache efficiency, a comprehensive strategy is required, factoring in each layer’s attention pattern and KV cache preference. Moreover, conventional methods that evict KV pairs relying solely on static spatial attention scores often overlook the temporal evolution of attention, missing critical insights into how token relevance changes in various contexts.
@@
Figure 2: Illustration of CAKE compared with existing cache allocation strategies. (a) Uniform cache allocation (Xiao et al., 2023; Zhang et al., 2024b; Li et al., 2024b); (b) Fixed-shape cache allocation (Cai et al., 2024; Yang et al., 2024a); (c) Preference-prioritized adaptive cache allocation used in CAKE. Compared to (a) and (b), CAKE adjusts allocation ratios across different layers with layer preferences, adapting to various contexts and models, with given memory budgets.
@@ 2 Background and Related Works
KV Cache Eviction. KV cache eviction during model inference enhances computational efficiency without altering the attention mechanism. This optimization relies on strategic cache reduction techniques. Early approaches to KV cache eviction focus on specific parts of the input sequence. For instance, StreamingLLM (Xiao et al., 2023) and LM-Infinite (Han et al., 2023) prioritize the retention of the first and last tokens. However, this strategy risks ignoring potentially important tokens in the middle of the sequence. Recognizing this limitation, subsequent research introduces more sophisticated indicators to filter unnecessary KV cache entries, such as using cumulative attention scores  (Zhang et al., 2024b; Liu et al., 2024d), last token attention scores (Oren et al., 2024), mean attention scores (Ren & Zhu, 2024), or clustering recent attention scores (Li et al., 2024b). While these methods offer more nuanced approaches to cache eviction, they often apply the uniform allocation strategy, which may not maximize cache utilization. Recent FastGen (Ge et al., 2023) chooses which tokens to keep based on special markers, but it doesn’t limit the total cache size. PyramidInfer (Yang et al., 2024a) and PyramidKV (Cai et al., 2024) allocate the budget in a pyramid-shaped manner, while D2O (Wan et al., 2024) adjusts cache size based on the current layer’s attention density. While these methods have shown promise, they often rely on predefined cache allocation strategies across all layers or are just based on the local layer attention, which does not fully capture the complex dynamics of attention mechanisms and lacks a global perspective on layer preferences for the cache. In contrast, our work addresses these limitations by considering layer-specific cache preferences with a comprehensive and global perspective.
Figure 3: Analysis of attention dynamics. (a) Heatmap for spatial attention dispersion (upper) and temporal attention shift (lower), red color with high value, while blue for low value. The x-axis represents samples, and the y-axis represents layers. (b) Variation of spatial attention dispersion (upper) and temporal attention shift (lower) across layers and models. The experimental data is derived from the LongBench dataset (Bai et al., 2023).
@@ 3 Insights into Attention Dynamics
Recent studies (Xiao et al., 2023; Zhang et al., 2024b) have shown that only a few KV cache elements are crucial during token decoding. Building on this, our analysis identifies two pivotal attention traits: spatial attention dispersion, showing how one token’s attention spreads across others, and temporal attention shift, tracking the evolution of highly-attended tokens over time.
@@
We analyze attention dynamics across multiple LLM layers using the LongBench dataset’s long-context samples (Bai et al., 2023). Figure 3 (upper) shows attention dispersion changes across layers, and Figure 3 (lower) shows attention focus shifts. The results show significant variations in attention dispersion and shift across layers, models, and contexts. Despite recent KV cache advances (Xiao et al., 2023; Zhang et al., 2024b), our findings highlight LLM attention complexity, indicating a need for tailored KV cache management strategies to effectively handle the dynamic nature of attention.
@@ 4.1 Preference-Prioritized Adaptive Allocation
where $\mathcal{P}$ is the preference score for an attention layer’s cache size. As dispersed attention (high $\mathcal{H}$) requires retaining broader contexts, while frequent shifts (high $\mathcal{V}$) demand support for complex temporal patterns, layers with high preference score $\mathcal{P}$ benefit more from larger KV cache to maintain performance. Accounting for the varying importance of attention dispersion and shift across models and memory constraints, we introduce temperature parameters $\tau_{1}$ and $\tau_{2}$ to adjust their influence on $\mathcal{P}$. We focus on the submatrix $\mathbf{A}[-S_{w}:,:-S_{w}]$ of $\mathbf{A}$, representing a recent window of size $S_{w}$. This focus on the recent window chunk of prefilling attention weights is inspired by recent research (Li et al., 2024b; Yang et al., 2024a). These studies have demonstrated that decoding patterns can be effectively captured by analyzing the attention patterns of recent queries.
@@
6:     $\mathbf{P}^{(m)}\leftarrow\mathbf{P}^{(m-1)}\cup(\mathcal{P}_{m})$,$\mathbf{S}^{(m)}\leftarrow\mathbf{S}^{(m-1)}\cup(\mathbf{I}_{m}^{(m)})$
@@ 4.2 Preference-Guided Cascading Cache Management
For any layer $l\in[L]$, the allocated budget size decreases monotonically from stage $l$ to $L-1$:
@@
For any layer $l\in[L]$, the KV cache $\{\mathbf{K}_{l},\mathbf{V}_{l}\}^{(L-1)}$ obtained through cascading eviction from stage $l$ to stage $L-1$ is equivalent to the result of applying the eviction operation once on the full KV cache $\{\mathbf{K}_{l},\mathbf{V}_{l}\}\in\mathbb{R}^{S\times D}$ with the cache budget $B_{l}$ computed in Eq.(7):
@@ 4.3 Attention-Shift Tolerant Eviction Indicator
Current top eviction indicators, such as accumulated or mean attention scores (Zhang et al., 2024b; Liu et al., 2024d; Ren & Zhu, 2024), simplify tokens’ attention profiles into single values, effectively identifying important tokens. However, this approach may overlook tokens whose importance fluctuates, as it fails to capture the dynamics of shifting attention. This could result in prematurely evicting tokens from the cache, affecting the model’s future ability to retrieve relevant information.
@@
where $\text{Mean}(\cdot)$ and $\text{Var}(\cdot)$ measure sustained importance and attention variability, respectively, with $\gamma$ adjusting their influence. Inspired by recent research (Li et al., 2024b; Yang et al., 2024a), we assign an arbitrarily large value $\Omega$ to ensure the preservation of the most recent $W$ tokens. Also, a pooling layer clusters $\mathbf{I}_{l}[:S-S_{w}]$ to maintain context and prevent information fragmentation.
@@ 5.1 Experimental Settings
Backbone LLMs. Our experiments involve five major open-source LLMs (7B-70B parameters) capable of handling 4k-128k token contexts. These models feature two representative attention structures: (1) Multi-head attention: including Llama2-Chat (Touvron et al., 2023) and Gemma-Instruct (Team et al., 2024). (2) Grouped-query attention: including Llama3-Instruct (Dubey et al., 2024), Mistral-v0.3 (Jiang et al., 2023), and Qwen2.5-Instruct (Team, 2024).
Baseline Methods. We assess CAKE’s performance against five baselines: (1) Uniform Allocation: StreamingLLM (Xiao et al., 2023) balances initial and recent tokens, H2O (Zhang et al., 2024b) uses cumulative attention, TOVA (Oren et al., 2024) adopts last-token attention, and SnapKV (Li et al., 2024b) clusters recent attention. (2) Non-Uniform Allocation: PyramidKV (Cai et al., 2024) employs a fixed pyramid-shaped allocation strategy with SnapKV’s eviction indicator.
Evaluating Tasks. To evaluate CAKE’s performance across various memory budgets, we use two carefully designed benchmarks: (1) LongBench (Bai et al., 2023): Focuses on long-context understanding, encompassing 16 datasets in six categories: Single/Multi-Document QA, Summarization, Few-shot Learning, Synthetic Tasks, and Code Completion. (2) NeedleBench (Li et al., 2024a): Tests retrieval and reasoning in complex contexts through three subtasks: Single-Needle Retrieval, Multi-Needle Retrieval, and Multi-Needle Reasoning.
@@ 5.4 Evaluation of Memory and Throughput
We assess the effectiveness of our method in reducing memory consumption and enhancing time efficiency during LLM inference by analyzing peak memory usage and decoding latency on Mistral-7B-Instruct-v0.3 implemented with FlashAttention-2 (Dao, 2023). Our comparison includes full cache, and three KV cache eviction methods, all with $B_{\text{total}}=1024L$: SnapKV (uniform allocation strategy), PyramidKV (fixed-pattern allocation strategy), and our proposed CAKE method.
@@
Throughput Analysis. Regarding decoding latency, the right panel of Figure 8 shows that both uniform and non-uniform eviction methods exhibit comparable inference performance. As input length increases, decoding latency for the full cache method grows significantly due to escalating computational demands and I/O latency bottlenecks. In contrast, CAKE maintains a relatively stable decoding speed by preserving a fixed amount of KV cache, resulting in significantly lower latency compared to the full cache, particularly for longer sequences. It is noteworthy that CAKE demonstrates remarkable efficiency, achieving over 10$\times$ speedup in decoding latency compared to the full cache approach when processing sequences with 128K context length.
@@ Acknowledgement
The paper is supported in part by the National Natural Science Foundation of China (No.62325109, U21B2013) and Ant Group.
@@ Appendix A Additional Related Works
Efficient Attention. Numerous strategies have been devised to tackle the high time and space demands of the attention layer. These strategies include low-rank techniques to transform the original attention mechanism into one with linear or near-linear complexity (Choromanski et al., 2020; Wang et al., 2020; Katharopoulos et al., 2020; Alberti et al., 2023), sparse attention to confine the attention to specific, predetermined or adaptively learned patterns (Child et al., 2019; Kitaev et al., 2020; Roy et al., 2021), and their integration for a more comprehensive solution (Beltagy et al., 2020; Zaheer et al., 2020; Chen et al., 2021; Ren et al., 2021). Another group minimizes KV cache size. This is achieved by either re-engineering the architecture of multi-head attention or by innovating the update process of the KV cache. Multi-Query Attention (MQA) (Shazeer, 2019) and Group-Query Attention (GQA) (Ainslie et al., 2023) reduce KV heads by sharing KV pairs. Multi-head Latent Attention (MLA) (Liu et al., 2024a) furthers efficiency by compressing keys and values into latent vectors, while YOCO (Sun et al., 2024) employs cross-attention to recycle the shared KV cache. Dynamic Memory Compression (DMC) (Nawrot et al., 2024) dynamically consolidates the KV cache during the training process, thereby optimizing resource utilization.
Orthogonal KV Cache Compression. Beyond the eviction of the KV cache, numerous strategies have been investigated for its compression. Specifically, quantization methods (Liu et al., 2024e; Yue et al., 2024; Kang et al., 2024) have been applied to condense the KV cache into a lower-bitwidth format. Moreover, a fusion of KV cache eviction with quantization (Yang et al., 2024b), pruning Xu et al. (2024), low-rank factorization (Dong et al., 2024), or cache merging (Liu et al., 2024b; Wan et al., 2024) has been explored to achieve a more compact form of the cache, preserving some of the discarded data. Additionally, there are initiatives to transfer the evicted KV cache to auxiliary memory resources for potential redeployment in computation (Xiao et al., 2024; Fu et al., 2024). It is crucial to acknowledge that these compression techniques are compatible with KV cache eviction and, when integrated, can amplify their individual benefits, leading to more robust solutions for easing the KV cache load.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Large language models (LLMs) (zhao2023survey; achiam2023gpt; dubey2024llama; anthropic2024claude3; mistral2024large) have enhanced their long-text processing capabilities, improving performance in multi-turn dialogues (chiang2023vicuna), document summarization (zhang2024benchmarking), question answering (kamalloo2023evaluating), and information retrieval (liu2024lost). New models such as GPT-4 (achiam2023gpt), Claude 3.5 (anthropic2024claude3), LLaMA 3.1 (dubey2024llama) and Mistral Large 2 (mistral2024large) have extended token processing capacities beyond 128K. Expanded contexts necessitate a linear increase in key-value (KV) cache size, resulting in heavier inference-time memory burdens. shazeer2019fast; ainslie2023gqa partially address this issue by merging key-value heads during the training phase. However, optimizing key-value cache without additional training is crucial for efficient inference of long contexts under memory constraints, particularly in typical deployment scenarios where the model structure is fixed.
One way to maintain a manageable KV cache size on the fly is to remove some KV pairs (xiao2023efficient; zhang2024h2o; li2024snapkv). The idea is to eliminate less important KV pairs based on certain rules. Although recent methods have enhanced pair selection for removal, they typically assign uniform cache sizes across layers, disregarding layer-specific requirements. This approach can impair performance under memory constraints. Recent research is trying to improve this by looking at how attention works in different layers (yang2024pyramidinfer; cai2024pyramidkv). These methods, which depend on prior observations, might not work well for all input types and models. Another idea is to adjust the cache size based on the current attention pattern (wan2024d2o). This can help, but it may not always optimize overall performance.
@@
Allocating optimal cache sizes for different layers with a fixed memory budget is akin to “cake-slicing.” The crux of this challenge is assessing each layer’s affinity for KV cache, which is mirrored in its attention mechanisms. Our research delves into these mechanisms, examining spatial and temporal aspects across layers. Spatially, attention distribution across tokens can differ markedly between layers. Some layers may spread attention broadly, while others concentrate on specific tokens, as depicted in Figure 1(a) and Figure 1(b). Temporally, attention hotspots in certain layers shift over time in Figure 1(c), whereas in others, they remain constant in Figure 1(d). The variability of attention pattern necessitates a dynamic approach to memory allocation that goes beyond uniform (zhang2024h2o; liu2024scissorhands; ren2024efficacy; li2024snapkv), depicted in Figure 2(a), or fixed-pattern methods (cai2024pyramidkv), drawn in Figure 2(b). They lack the flexibility to adapt to the nuanced attention dynamics. To enhance cache efficiency, a comprehensive strategy is required, factoring in each layer’s attention pattern and KV cache preference. Moreover, conventional methods that evict KV pairs relying solely on static spatial attention scores often overlook the temporal evolution of attention, missing critical insights into how token relevance changes in various contexts.
@@
Figure 2: Illustration of CAKE compared with existing cache allocation strategies. (a) Uniform cache allocation (xiao2023efficient; zhang2024h2o; li2024snapkv); (b) Fixed-shape cache allocation (cai2024pyramidkv; yang2024pyramidinfer); (c) Preference-prioritized adaptive cache allocation used in CAKE. Compared to (a) and (b), CAKE adjusts allocation ratios across different layers with layer preferences, adapting to various contexts and models, with given memory budgets.
@@ 2 Background and Related Works
KV Cache Eviction. KV cache eviction during model inference enhances computational efficiency without altering the attention mechanism. This optimization relies on strategic cache reduction techniques. Early approaches to KV cache eviction focus on specific parts of the input sequence. For instance, StreamingLLM (xiao2023efficient) and LM-Infinite (han2023lm) prioritize the retention of the first and last tokens. However, this strategy risks ignoring potentially important tokens in the middle of the sequence. Recognizing this limitation, subsequent research introduces more sophisticated indicators to filter unnecessary KV cache entries, such as using cumulative attention scores  (zhang2024h2o; liu2024scissorhands), last token attention scores (oren2024transformers), mean attention scores (ren2024efficacy), or clustering recent attention scores (li2024snapkv). While these methods offer more nuanced approaches to cache eviction, they often apply the uniform allocation strategy, which may not maximize cache utilization. Recent FastGen (ge2023model) chooses which tokens to keep based on special markers, but it doesn’t limit the total cache size. PyramidInfer (yang2024pyramidinfer) and PyramidKV (cai2024pyramidkv) allocate the budget in a pyramid-shaped manner, while D2O (wan2024d2o) adjusts cache size based on the current layer’s attention density. While these methods have shown promise, they often rely on predefined cache allocation strategies across all layers or are just based on the local layer attention, which does not fully capture the complex dynamics of attention mechanisms and lacks a global perspective on layer preferences for the cache. In contrast, our work addresses these limitations by considering layer-specific cache preferences with a comprehensive and global perspective.
Figure 3: Analysis of attention dynamics. (a) Heatmap for spatial attention dispersion (upper) and temporal attention shift (lower), red color with high value, while blue for low value. The x-axis represents samples, and the y-axis represents layers. (b) Variation of spatial attention dispersion (upper) and temporal attention shift (lower) across layers and models. The experimental data is derived from the LongBench dataset (bai2023longbench).
@@ 3 Insights into Attention Dynamics
Recent studies (xiao2023efficient; zhang2024h2o) have shown that only a few KV cache elements are crucial during token decoding. Building on this, our analysis identifies two pivotal attention traits: spatial attention dispersion, showing how one token’s attention spreads across others, and temporal attention shift, tracking the evolution of highly-attended tokens over time.
@@
We analyze attention dynamics across multiple LLM layers using the LongBench dataset’s long-context samples (bai2023longbench). Figure 3 (upper) shows attention dispersion changes across layers, and Figure 3 (lower) shows attention focus shifts. The results show significant variations in attention dispersion and shift across layers, models, and contexts. Despite recent KV cache advances (xiao2023efficient; zhang2024h2o), our findings highlight LLM attention complexity, indicating a need for tailored KV cache management strategies to effectively handle the dynamic nature of attention.
@@ 4.1 Preference-Prioritized Adaptive Allocation
where $\mathcal{P}$ is the preference score for an attention layer’s cache size. As dispersed attention (high $\mathcal{H}$) requires retaining broader contexts, while frequent shifts (high $\mathcal{V}$) demand support for complex temporal patterns, layers with high preference score $\mathcal{P}$ benefit more from larger KV cache to maintain performance. Accounting for the varying importance of attention dispersion and shift across models and memory constraints, we introduce temperature parameters $\tau_{1}$ and $\tau_{2}$ to adjust their influence on $\mathcal{P}$. We focus on the submatrix $\mathbf{A}[-S_{w}:,:-S_{w}]$ of $\mathbf{A}$, representing a recent window of size $S_{w}$. This focus on the recent window chunk of prefilling attention weights is inspired by recent research (li2024snapkv; yang2024pyramidinfer). These studies have demonstrated that decoding patterns can be effectively captured by analyzing the attention patterns of recent queries.
@@
6:  $\mathbf{P}^{(m)}\leftarrow\mathbf{P}^{(m-1)}\cup(\mathcal{P}_{m})$$\mathbf{S}^{(m)}\leftarrow\mathbf{S}^{(m-1)}\cup(\mathbf{I}_{m}^{(m)})$
@@ 4.2 Preference-Guided Cascading Cache Management
For any layer $l\in[L]$, the allocated budget size decreases monotonically from stage $l$ to L1L-1:
@@
For any layer $l\in[L]$, the KV cache $\{\mathbf{K}_{l},\mathbf{V}_{l}\}^{(L-1)}$ obtained through cascading eviction from stage $l$ to stage L1L-1 is equivalent to the result of applying the eviction operation once on the full KV cache $\{\mathbf{K}_{l},\mathbf{V}_{l}\}\in\mathbb{R}^{S\times D}$ with the cache budget $B_{l}$ computed in Eq.(7):
@@ 4.3 Attention-Shift Tolerant Eviction Indicator
Current top eviction indicators, such as accumulated or mean attention scores (zhang2024h2o; liu2024scissorhands; ren2024efficacy), simplify tokens’ attention profiles into single values, effectively identifying important tokens. However, this approach may overlook tokens whose importance fluctuates, as it fails to capture the dynamics of shifting attention. This could result in prematurely evicting tokens from the cache, affecting the model’s future ability to retrieve relevant information.
@@
where $\text{Mean}(\cdot)$ and $\text{Var}(\cdot)$ measure sustained importance and attention variability, respectively, with $\gamma$ adjusting their influence. Inspired by recent research (li2024snapkv; yang2024pyramidinfer), we assign an arbitrarily large value $\Omega$ to ensure the preservation of the most recent $W$ tokens. Also, a pooling layer clusters $\mathbf{I}_{l}[:S-S_{w}]$ to maintain context and prevent information fragmentation.
@@ 5.1 Experimental Settings
Backbone LLMs. Our experiments involve five major open-source LLMs (7B-70B parameters) capable of handling 4k-128k token contexts. These models feature two representative attention structures: (1) Multi-head attention: including Llama2-Chat (touvron2023llama) and Gemma-Instruct (team2024gemma). (2) Grouped-query attention: including Llama3-Instruct (dubey2024llama), Mistral-v0.3 (jiang2023mistral), and Qwen2.5-Instruct (qwen2.5).
Baseline Methods. We assess CAKE’s performance against five baselines: (1) Uniform Allocation: StreamingLLM (xiao2023efficient) balances initial and recent tokens, H2O (zhang2024h2o) uses cumulative attention, TOVA (oren2024transformers) adopts last-token attention, and SnapKV (li2024snapkv) clusters recent attention. (2) Non-Uniform Allocation: PyramidKV (cai2024pyramidkv) employs a fixed pyramid-shaped allocation strategy with SnapKV’s eviction indicator.
Evaluating Tasks. To evaluate CAKE’s performance across various memory budgets, we use two carefully designed benchmarks: (1) LongBench (bai2023longbench): Focuses on long-context understanding, encompassing 16 datasets in six categories: Single/Multi-Document QA, Summarization, Few-shot Learning, Synthetic Tasks, and Code Completion. (2) NeedleBench (li2024needlebench): Tests retrieval and reasoning in complex contexts through three subtasks: Single-Needle Retrieval, Multi-Needle Retrieval, and Multi-Needle Reasoning.
@@ 5.4 Evaluation of Memory and Throughput
We assess the effectiveness of our method in reducing memory consumption and enhancing time efficiency during LLM inference by analyzing peak memory usage and decoding latency on Mistral-7B-Instruct-v0.3 implemented with FlashAttention-2 (dao2023flashattention2fasterattentionbetter). Our comparison includes full cache, and three KV cache eviction methods, all with $B_{\text{total}}=1024L$: SnapKV (uniform allocation strategy), PyramidKV (fixed-pattern allocation strategy), and our proposed CAKE method.
@@
[…]
@@ Acknowledgement
The paper is supported in part by Ant Group through the CCF-Ant Research Fund and the National Natural Science Foundation of China (No.62325109, U21B2013).
@@ Appendix A Additional Related Works
Efficient Attention. Numerous strategies have been devised to tackle the high time and space demands of the attention layer. These strategies include low-rank techniques to transform the original attention mechanism into one with linear or near-linear complexity (choromanski2020rethinking; wang2020linformer; katharopoulos2020transformers; alberti2023sumformer), sparse attention to confine the attention to specific, predetermined or adaptively learned patterns (child2019generating; kitaev2020reformer; roy2021efficient), and their integration for a more comprehensive solution (beltagy2020longformer; zaheer2020big; chen2021scatterbrain; ren2021combiner). Another group minimizes KV cache size. This is achieved by either re-engineering the architecture of multi-head attention or by innovating the update process of the KV cache. Multi-Query Attention (MQA) (shazeer2019fast) and Group-Query Attention (GQA) (ainslie2023gqa) reduce KV heads by sharing KV pairs. Multi-head Latent Attention (MLA) (liu2024deepseek) furthers efficiency by compressing keys and values into latent vectors, while YOCO (sun2024you) employs cross-attention to recycle the shared KV cache. Dynamic Memory Compression (DMC) (nawrot2024dynamic) dynamically consolidates the KV cache during the training process, thereby optimizing resource utilization.
Orthogonal KV Cache Compression. Beyond the eviction of the KV cache, numerous strategies have been investigated for its compression. Specifically, quantization methods (liu2024kivi; yue2024wkvquant; kang2024gear) have been applied to condense the KV cache into a lower-bitwidth format. Moreover, a fusion of KV cache eviction with quantization (yang2024no), pruning xu2024think, low-rank factorization (dong2024get), or cache merging (liu2024minicache; wan2024d2o) has been explored to achieve a more compact form of the cache, preserving some of the discarded data. Additionally, there are initiatives to transfer the evicted KV cache to auxiliary memory resources for potential redeployment in computation (xiao2024infllm; fu2024lazyllm). It is crucial to acknowledge that these compression techniques are compatible with KV cache eviction and, when integrated, can amplify their individual benefits, leading to more robust solutions for easing the KV cache load.